Monitors manufactured by TSI Electronics have life spans that have a normal distribution with a variance of 2,250,000 and a mean life span of 13,000 hours. If a monitor is selected at random, find the probability that the life span of the monitor will be more than 14,650 hours. Round your answer to four decimal places.
Answer #1

Solution :

Given that ,

mean = = 13000 hours

variance = 2 = 2,250,000 hours

standard deviation = = 2 = 2,250,000 = 1500 hours

P(x > 14650) = 1 - p( x< 14650)

=1- p P[(x - ) / < (14650 - 13000) / 1500 ]

=1- P(z < 1.10)

Using z table,

= 1 - 0.8643

= 0.1357